### v1

# Comparing Neighbors Together Makes it Easy: Jointly Comparing Multiple Candidates for Efficient and Effective Retrieval

Abstract
A common retrieve-and-rerank paradigm involves retrieving a broad set of relevant candidates using a scalable bi-encoder, followed by expensive but more accurate cross-encoders to a limited candidate set. However, this small subset often leads to error propagation from the bi-encoders, thereby restricting the performance of the overall pipeline. To address these issues, we propose the Comparing Multiple Candidates (CMC) framework, which compares a query and multiple candidate embeddings jointly through shallow self-attention layers. While providing contextualized representations, CMC is scalable enough to handle multiple comparisons simultaneously, where comparing 2K candidates takes only twice as long as comparing 100. Practitioners can use CMC as a lightweight and effective reranker to improve top-1 accuracy. Moreover, when integrated with another retriever, CMC reranking can function as a virtually enhanced retriever. This configuration adds only negligible latency compared to using a single retriever (virtual), while significantly improving recall at K (enhanced). Through experiments, we demonstrate that CMC  as a virtually enhanced retriever, significantly improves Recall@k (+6.7, +3.5%-p for R@16, R@64) compared to the initial retrieval stage on the ZeSHEL dataset. Meanwhile, we conduct experiments for direct reranking on entity, passage, and dialogue ranking. The results indicate that CMC is not only faster (11x) than cross-encoders but also often more effective, with improved prediction performance in Wikipedia entity linking (+0.7%-p) and DSTC7 dialogue ranking (+3.3%-p). The code and link to datasets are available at https://anonymous.4open.science/r/cmd/
Comparing Neighbors Together Makes it Easy: Jointly Comparing Multiple Candidates for Efficient and Effective Retrieval
1 Introduction
Figure 1: Overview of ranking model architectures for sequence retrieval tasks. Model (a), (b), and (c) are existing architectures. Model (d) is our proposed ‘Comparing Multiple Candidates (CMC)’ architecture, which computes compatibility scores by comparing the embeddings of a query and K multiple candidates via self-attention layers. Contarty to (a)-(c), our model can process multiple documents at once rather than conducting multiple forward passes of (query, candidate) pair for each document.
The two-stage approach of retrieval and reranking has become a predominant approach for open-domain question answering (ODQA) Nogueira and Cho (2019); Agarwal et al. (2022b); Shen et al. (2022); Qu et al. (2020), entity linking (EL) Wu et al. (2020); Zhang and Stratos (2021); Xu et al. (2023), and dialogue systems Mele et al. (2020). Typically, bi-encoders (BE) are used to efficiently retrieve relevant candidates among a large set of documents (e.g. knowledge base), and then cross-encoders (CE) effectively rerank only a confident subset of candidates already retrieved by BE Nogueira and Cho (2019) (Figure 1.a-b).
The two-stage approach, although widely used, has limitations: it has an innate trade-off between efficiency and effectiveness and is susceptible to error propagation. When inaccurate BE retrieve too few candidates, the framework risks missing the gold candidates. Once the ground truth is not retrieved, no subsequent reranking can consider the gold candidates anymore, leading to the failure of the entire system. In contrast, simply increasing the number of candidates introduces another problem of slow processing time that limits scalability under time constraints. Furthermore, increasing the number of candidates for CE does not necessarily improve end-to-end accuracy  [^1] We also confirm this in the experiments. See appendix E.6. Wu et al. (2020). Therefore, the retrieval framework poses a dilemma for users between the inaccurate performance of BE and the slow runtime of CE.
To resolve this problem, several studies have explored various strategies to find an optimal balance in the efficiency-effectiveness tradeoff. Khattab and Zaharia (2020); Zhang and Stratos (2021); Cao et al. (2020); Humeau et al. (2019) have enhanced bi-encoder architectures by adding an interaction component between the query and the candidate. However, these late interaction models require to save multiple vector embeddings for each candidate in advance, or only utilize single query-candidate interaction without considering the relationships between candidates (Figrue 1.c).
In this paper, we present a novel reranking framework CMC, which stands for Comparing Multiple Candidates (Figure 1.d). CMC is designed to improve both accuracy and scalability by comparing multiple candidates along with the query simultaneously. It compares a query and multiple pre-computed candidate embeddings jointly through a few layers of bi-directional self-attention layers. In other words, CMC only takes one forward pass for single input $(\text{query},\text{candidate}_{1},...,\text{candidate}_{k})$, while CE and other late interaction models take $k$ separate forward passes for multiple input pairs $(\text{query},\text{candidate}_{1}),...,(\text{query},\text{candidate}_{k})$. The proposed configuration of CMC  maintains both the efficiency of BE by using pre-computed candidate embeddings, and the effectiveness of CE by using attention mechanisms between query and multiple candidates (Figure 1.d).
Practitioners can use CMC as a fast and effective reranker, capable of enhancing top-1 retrieval (↑R@1). Also, its efficiency enables CMC to function as a virtually enhanced retrieval system. When integrated with another retriever, such as BE, CMC efficiently identifies better candidates (enhanced; ↑R@K) from a large pool with minimal additional latency (vitrual; Figure 2). As slow cross-encoders can only process a limited number of candidates, filtering higher-quality candidates with CMC contributes to minimizing the error propagation in the retrieval process.
In experiments, we evaluate CMC on Zero-SHot Entity-Linking dataset (ZeSHEL; Logeswaran et al. (2019)) to investigate the performance as a virtually enhanced retriever. The results show that CMC provides higher recall than existing state-of-the-art retrievers (+3.5 Recall@64 points) at marginal computing costs (+0.07x runtime of BE; Table 5). Additionally, we demonstrate that CMC provides high-quality candidates that enhance CE’s end-to-end prediction accuracy even with a reduced candidate set, while improving overall latency (Table 2). To examine the effectiveness of CMC as a reranker, we also evaluate CMC on entity, passage, and dialogue ranking tasks. We observe that CMC outperforms CE on both Wikipedia entity linking datasets (+0.7 accuracy points) and DSTC7 dialogue ranking datasets (+3.3 MRR points), requiring only 0.09x of CE’s reranking latency (Table 3).
The main contribution of the paper is as follows:
  • We propose a novel CMC that is both effective and scalable reranker by utilizing single pre-computed candidate embeddings (§3).
  • We show that CMC can function as a virtually enhanced retriever, which increases the recall of the initial retriever at a marginal cost and thus improves overall reranking performance. (§4.3)
  • We provide experimental results which show CMC has robust performance on passage ranking, entity linking, and dialogue ranking tasks compared to various baselines (§4.4).
  • Additionally, we show that CMC can benefit from domain transfer from sentence encoders while BE cannot (§4.5).
Figure 2: Overview of the proposed CMC framework that compares multiple candidates at once. The configuration of Bi-encoder + CMC can be used as both a virtually enhanced retriever which finds top-K’ candidates or a reranker which outputs top-1 candidate. Candidate embeddings for bi-encoders and CMC are both precomputed while query embeddings for bi-encoders and CMC are computed in parallel on the fly. After bi-encoders retrieve top-$K$ candidates, CMC indexes the corresponding candidate embeddings and passes through a two-layer transformer encoder.
2 Background and Related Works
2.1 Retrieve and rerank
Information retrieval systems commonly comprise two stages: efficient retrieval and effective reranking. A fast retriever will score the query $q$ with each candidate $c\in\mathcal{C}$. Although the retriever is designed to be fast, its top-1 accuracy tends to be suboptimal. Practitioners therefore identify a candidate set.
$C_{q}=\{c_{q,1},c_{q,2},\ldots,c_{q,K}\}\subseteq\mathcal{C},$ (1)
whose elements are $K$ most relevant candidates in corpus $\mathcal{C}$, to be reranked.
A reranker $s_{\theta}(q,c_{q,j})(1\leq j\leq K)$ is a model learned to assign more fine-grained matching scores between a query $q$ and each candidate $c_{q,j}$ from the relatively small set of candidates $C_{q}$ predicted by the retriever. It is an expressive model that is generally slower but more accurate than the retriever. The candidate with the highest score,
$\hat{c}_{q}=\operatorname*{arg\,max}_{c_{q,j}\in C_{q}}s_{\theta}(q,c_{q,j}),$ (2)
is the final output where query $q$ should be linked.
2.2 Related Work
Bi-encoders and Cross-encoders
In the two-stage retrieval system, the compatibility score between the query and candidate can be computed by diverse functions. Nogueira et al. (2019a) first retrieves candidates for a query using the bag-of-words BM25 retriever and then uses a cross-encoders, transformer encoders that input the query and candidate tokens jointly, to get a final score. Numerous works Lee et al. (2019); Gillick et al. (2019); Karpukhin et al. (2020) use a pre-trained language model for bi-encoders architecture to encode a query and a candidate separately and get the compatibility score between the them. The scalability of bi-encoders comes from the indexing of candidates and maximum inner-product search (MIPS); however, they are less effective than cross-encoders as candidate representation does not reflect query information (Figure 1 (a-b)). To enhance the performance of bi-encoders, follow-up works propose a task-specific fine-tuned model Gao and Callan (2022), injecting graph information Wu et al. (2023); Agarwal et al. (2022a), and multi-view text representation Ma et al. (2021); Liu et al. (2023).
Late Interaction
Late interaction models, which typically function as either a retriever or a reranker, enhance bi-encoder architectures with an interaction component between the query and candidates.
Poly-encoder Humeau et al. (2019) and MixEncoder Yang et al. (2023) represent query information through cross-attention with individual candidates to calculate matching scores. However, these models have overlooked the opportunity to explore the interaction across candidates.
Sum-of-Max Khattab and Zaharia (2020); Zhang and Stratos (2021); Santhanam et al. (2021) and DeFormer Cao et al. (2020) rely on Maximum Similarity (MaxSim) operations or additional cross-encoder layers on top of bi-encoders. These approaches increase offline indexing costs as they require every token embedding for each candidate, complicating their application to tasks where the knowledge base is frequently updated and the index must be continuously revised.
CMC diverges from these models by using a single vector embedding for each candidate, enabling interactions among multiple candidates with enhanced scalability. It helps to explore deeper relational dynamics among candidates while enhancing memory efficiency.
Listwise Ranking
CMC is not the first approach to jointly compare a query and documents to enhance ranking performance. This listwise ranking method has been explored in previous works Han et al. (2020); Zhang et al. (2022); Xu et al. (2023). They process cross-encoder logits for the list $(\text{query},\text{candidate}_{1},\ldots,\text{candidate}_{\text{K}})$ to rerank $K$ candidates from cross-encoders. Although they focus on performance, these approaches lack scalability as depending on pairwise representations from cross-encoders.
Unlike previous listwise ranking models with pairwise representations from cross-encoders, we propose a method that employs single representations from sentence encoders. In particular, CMC can function as a virtually enhanced retriever, boosting scalability by using independent sentence encoders for query and candidate, while still maintaining accurate predictions.
3 Proposed Method
3.1 Model Architecture
Comparing Multiple Candidates, CMC, employs self-attention layers to capture both query-candidate and candidate-candidate interactions. Contrary to other late interaction models Khattab and Zaharia (2020); Humeau et al. (2019); Yang et al. (2023), which compute the matching scores by only comparing a query and an individual candidate, CMC  considers the query in conjunction with all candidates simultaneously. In other words, CMC enriches each candidate’s representation by contextualizing it with other candidates and the query. In addition, this architecture enables the processing of multiple candidates in a scalable manner by pre-computing individual candidate embeddings as discussed in §3.2. The overall architecture in Figure 1.(d) provides a high-level picture of the CMC, describing how it integrates and processes multiple candidate interactions given a single query. In detail, the self-attention layer is designed to process the aggregated encoder output (i.e. $[CLS]$ token embedding) of the query and multiple candidates. The outputs are derived from our query encoder $\texttt{Enc}_{qry}$ and candidate encoder $\texttt{Enc}_{can}$.
Query and Candidate Encoders
Prior to CMC, the initial retriever identifies the candidate set with K elements $C_{q}=\{c_{q,1},...c_{q,K}\}$ for query $q$. Each query and candidate is represented by tokens $\text{x}_{q}$ and $\text{x}_{c_{q,j}}$. Initially, CMC obtains the aggregated output of query sentence tokens ($\mathbf{h}^{sent}_{q}$) and candidate sentence tokens ($\mathbf{h}^{sent}_{c_{q,j}}$) from the query encoder $\texttt{Enc}_{qry}$ and the candidate encoder $\texttt{Enc}_{can}$. These encoders play the same role as conventional bi-encoders in that condensing each query and candidate information into single vector embedding, but are trained separately from bi-encoders for the initial stage retriever.
$\mathbf{h}^{sent}_{q}$ $=\texttt{aggregator}(\texttt{Enc}_{qry}([\texttt{CLS}]\text{x}^{0}_{q}\ldots\text{x}^{k}_{q}))$ (3)
$\mathbf{h}^{sent}_{c_{q,j}}$ $=\texttt{aggregator}(\texttt{Enc}_{can}([\texttt{CLS}]\text{x}^{0}_{c_{q,j}}\ldots\text{x}^{k}_{c_{q,j}}))$ (4)
The aggregator function extracts [CLS] token embeddings from the last layer of each encoder [^2] For entity linking tasks, both the query (mention) and candidate (entity) sentences include custom special tokens that denote the locations of mention and entity words. These include [SEP], [query_start], [query_end], and [DOC] tokens following Wu et al. (2020)..
Self-attention Layer
The self-attention layer jointly processes concatenated embeddings of a query and all candidate sentences. This lightweight module enables parallel computation (efficient) and generates contextualized embeddings that capture query-candidate and candidate-candidate interaction (effective). Representing candidates together with self-attention enables fine-grained comparison among candidates. The self-attention layer consists of two layers of vanilla transformer encoder Vaswani et al. (2017) in Pytorch without positional encoding.
$\begin{split}\left[\mathbf{h}^{\texttt{CMC}}_{q};\mathbf{h}^{\texttt{CMC}}_{c_{q,1}};\ldots;\mathbf{h}^{\texttt{CMC}}_{c_{q,K}}\right]\\ =\texttt{SelfAttn}\Big{(}&\left[\mathbf{h}^{sent}_{q};\mathbf{h}^{sent}_{c_{q,1}};\ldots;\mathbf{h}^{sent}_{c_{q,K}}\right]\Big{)}\end{split}$ (5)
Consequently, CMC computes the final prediction $\hat{c}_{q}$ via dot products of query and candidate embeddings from the self-attention layer:
$\hat{c}_{q}=\operatorname*{arg\,max}_{c_{q,j}\in C_{q}}\mathbf{h}^{\texttt{CMC}}_{q}\cdot\left(\mathbf{h}^{\texttt{CMC}}_{c_{q,j}}\right)^{\top}$ (6)
3.2 Inference
Offline Indexing
CMC is capable of pre-computing and storing candidate embeddings offline, unlike cross-encoders Logeswaran et al. (2019); Wu et al. (2020) (Figure 1). CMC processes the candidates in the collection (e.g. knowledge base) and stores a single embedding for each. As demonstrated in Section 4.4, offline indexing significantly reduces the inference time compared to that of cross-encoders. Furthermore, it allows the runtime performance of CMC to be comparable to that of bi-encoders. While reducing time complexity, the space requirement for our model is less than $1\%$ of that required by Sum-of-Max and Deformer, as these models necessitate every token embedding for scoring, whereas CMC requires only a single vector embedding per candidate.
Parallel Computation
Given multiple GPUs, the end-to-end runtime for retrieving and reranking with CMC  can be comparable to that of bi-encoders retrieval. This is achieved through parallel computation of query encoders at bi-encoders and CMC (Figure 2). Consequently, the additional latency for running CMC is limited to the execution of a few encoder stacks within the self-attention layer.
3.3 CMC as a Virtually Enhanced Retriever
The parallel computation of CMC (Section 3.2) offers an experience of virtually enhanced retriever, requiring only a slight increase in computing resources (virtual) while significantly improving recall at various K (↑R@K) (enhanced). This process begins with initial retrievers such as bi-encoders, which retrieve a broad set of candidates. CMC then narrows this set down to a more manageable number for cross-encoders (e.g., 64 or fewer). Since CMC can rerank candidates from the initial retriever with only a marginal increase in latency, the runtime for bi-encoders is comparable to that of a virtually enhanced retriever running bi-encoders and CMC in parallel. Consequently, the improved recall contributes to the performance increase of the final stage reranker (e.g., cross-encoders) at a marginal cost, optimizing overall performance with comparable efficiency (Section 4.3).
3.4 CMC as a Reranker
The obvious application of CMC is the final stage reranker whose goal is to increase R@1 as well as R@K. Effective reranking is achieved by contextualizing the query and the candidates concerning each other while maintaining efficiency by utilizing a single vector embedding for each. In the training stage, CMC is typically fed with 64 candidates per query. Surprisingly, this training setup remains effective for a varying number of candidates at the inference stage. For example, even though CMC was trained with 64 candidates on the MS MARCO passage ranking dataset, it shows effectiveness when handling up to 1,000 candidates (Section §4.4). This evidence not only emphasizes the scalability of CMC but also its robustness in processing a diverse distribution of candidates.
4 Experiments
4.1 Dataset
To assess the empirical impact of CMC , we conduct our experiments on entity linking, passage ranking, and dialogue ranking tasks, where the retrieve-and-rerank approach is commonly employed. For entity linking, we utilize datasets linked to the Wikipedia knowledge base (AIDA-CoNLL Hoffart et al. (2011), WNED-CWEB Guo and Barbosa (2018), and MSNBC Cucerzan (2007)), as well as a ZEro-SHot Entity Linking dataset (ZeSHEL; Logeswaran et al. (2019)) based on the Wikia333now Fandom: https://www.fandom.com knowledge base. The candidates are retrieved from bi-encoders fine-tuned for each knowledge base Wu et al. (2020); Yadav et al. (2022). For passage ranking, we conduct an experiment on MS MARCO with 1K candidates from BM25 as an initial retriever following Bajaj et al. (2016). For dialogue ranking tasks, we test our model on DSTC7 challenge (Track 1) Yoshino et al. (2019), where conversations are extracted from technical support chat. For more details, see Appendix §C.
Test Speed Index Size
Method R@1 R@4 R@8 R@16 R@32 R@64 (ms) (GB)
Single- BM25 25.92 44.86 52.05 58.21 63.76 69.13
View Bi-encoder (BE) 52.94 64.51 71.94 81.52 84.98 87.95 568.9 0.2
Poly-encoder 40.04 60.16 67.23 72.16 76.49 80.18 581.0 0.2
Arbo-EL 50.31 68.34 74.26 78.4 82.02 85.11 - -
GER 42.86 66.48 73.00 78.11 81.11 85.65 - -
Sum-of-max 27.11 64.11 73.21 79.59 84.14 87.95 6393.0 25.7
BE + CMC(Ours) 59.08 77.58 82.91 86.32 89.33 91.51 607.2 0.4
Multi- MuVER 43.49 68.78 75.87 77.65 85.86 89.52 - -
View MVD 52.51 73.43 79.74 84.35 88.17 91.55 - -
MVD + CMC(Ours) 59.04 77.77 83.06 86.71 89.91 92.36 - -
Table 1: Retrieval performance over ZeSHEL dataset. The best result is denoted in bold and the second-best result is underlined for each method. BE is bi-encoders from Yadav et al. (2022) which is used for CMC. indicates our implementation as recall@k for all k are not provided in previous work555recall@64 of Poly-encoder and Sum-of-max from Zhang and Stratos (2021) is reported as 84.34 and 89.62, respectively.. Our result is conducted over top-512 candidates from the initial retriever and averaged over experiments with five random seeds.
4.2 Training Details
All the training details for CMC and other baselines are the same, following prior studies Wu et al. (2020); Zhang and Stratos (2021); Gao and Callan (2022); Yang et al. (2023). CMC and other baselines are trained using the AdamW optimizer with a 10% warmup step. Also, we investigate different sentence encoders for CMC and late interaction models: (1) vanilla BERT and (2) BERT-based sentence encoders fine-tuned on each dataset. After training, we select the best result among them. [^6] If more favorable results are found from prior works over the same candidates, we source the results from these works. For ZeSHEL, training CMC and other low-latency baselines for one epoch on a single NVIDIA A100 GPU takes approximately 4 hours. Every training detail for each dataset is outlined in Appendix §D.
4.3 CMC as a Virtually Enhanced Retriever
We conduct experiments on the ZeSHEL dataset to verify the impact of CMC as a virtually enhanced retriever. We conduct experiments on how the parallel computation of CMC reranking can achieve a higher recall@k with only a marginal increase in computational cost compared to conventional bi-encoders. We also evaluate how improved recall@k can enhance the performance of cross-encoders’ reranking. To do so, we compare how well cross-encoders rerank candidates filtered only by the bi-encoders versus those further refined by CMC.
Baselines
To assess CMC’s performance as a retriever, we compare CMC against baselines categorized into two types: single-view and multi-view retriever.  [^7] Single-view retrievers consider only a single global view derived from the entire sentence, whereas multi-view retrievers divide candidate information into multiple local views. As an initial retriever that provides candidates for CMC, we use bi-encoders fine-tuned for ZeSHEL Yadav et al. (2022) for the single-view retriever, and MVD Liu et al. (2023) for multi-view retriever. As baselines, we select baselines such as the poly-encoder Humeau et al. (2019), Sum-of-max Zhang and Stratos (2021), Arbo-EL Agarwal et al. (2022b), and GER Wu et al. (2023) for single-view retriever, and we include MuVER Ma et al. (2021) and MVD Liu et al. (2023), recognized as the SOTA retrievers for the ZeSHEL dataset.
Experimental Results
In Table 5, adopting CMC upon single-view retriever surpasses other baseline models across all k values, demonstrating its effectiveness in enhancing the end-to-end retrieval process. Also, CMC boosts recall@64 to 91.51% with candidates from the initial retriever whose recall@64 is 87.95%, with a marginal increase in latency (+0.07x). Similarly, operating CMC with MVD shows that virtually enhanced retriever consistently improves recall performance, regardless of the initial retriever used. It leads to an increase in performance, moving from 91.55 to 92.36 at recall@64. For the effect of the number of candidates from the initial retriever, see appendix §E.2.
To assess the effect of candidates retrieved from CMC on the overall performance, we analyze the subsequent cross-encoder’s reranking performance in Table 2. We explore various combinations where candidates are selected from either bi-encoders or CMC, in groups of (8, 16, 64). These candidates from CMC are initially selected from pools of (64, 256, 512) provided by bi-encoders which are the same or larger than 64 that typical cross-encoders consider.
A key finding presented in the table is that CMC’s scalability to search over a broad set enables it to provide better candidates for cross-encoders. By filtering candidates from bi-encoders to fewer, CMC effectively narrows down the most promising candidates for cross-encoders. In Table 2, the performances of cross-encoders with 16 candidates from CMC which filters out 256 or 512 candidates (row 8-9) are slightly better than their best results with 64 bi-encoder candidates (row 3), with a significant reduction in latency, making it 1.75 times faster by presenting fewer candidates for cross-encoders. Furthermore, cross-encoders reach their best performance with candidates from CMC. When using 64 candidates from CMC (row 11), cross-encoders surpass their performance with an equal number of bi-encoder candidates (row 3) by 0.17 points in macro accuracy.
Retrieved (k) Recall@k Unnormalized Accuracy Comparative
Row Bi-encoder CMC
Forgotten
Realms
Lego
Start
Trek
Yugioh
Macro
Avg.
Latency
1 8 - 77.72 78.92 65.14 62.76 48.64 63.87 38.90%
2 16 - 81.52 80.17 66.14 63.69 49.64 64.91 48.85%
3 64 - 87.95 80.83 67.81 64.23 50.62 65.87 100%
4 64 8 82.45 80.67 66.56 64.54 50.71 65.62 43.04%
5 256 8 82.86 80.92 66.89 64.42 50.86 65.77 43.36%
6 512 8 82.91 80.75 67.14 64.35 51.01 65.81 43.55%
7 64 16 85.46 80.5 66.97 64.47 50.68 65.66 56.76%
8 256 16 86.22 80.75 67.31 64.63 51.1 65.95 57.08%
9 512 16 86.22 80.83 67.64 64.49 50.95 65.98 57.27%
10 256 64 90.91 81.17 67.64 64.37 50.92 66.03 104.46%
11 512 64 91.51 81.00 67.89 64.42 50.86 66.04 104.65%
Table 2: Unnormalized accuracy of cross-encoders across various candidate configurations on the ZeSHEL dataset. We underlined when the cross-encoders show superior accuracy with candidates generated by CMC compared to those from bi-encoders. The top-performing scenarios in each category are highlighted in bold. We measure the comparative latency required for running cross-encoders over 64 bi-encoder candidates (260.84ms). For your reference, the CMC runtime 2x when increasing the number of candidates by 16x (from 128 to 1048), while able to compare up to 16k candidates at once. (Appendix E.1)
Tasks Entity Linking Passage Ranking Dialogue Ranking Compuational Efficiency
Datasets Wikipedia ZeSHEL MS MARCO Dev DSTC7 Challenge Total Speed Extra Memory
Accuracy Accuracy R@1 MRR@10 R@1 MRR@10
High-latency Cross-encoder 80.2±$\pm 0.2$ 65.9 25.4 36.8 64.7 73.2 12.9x -
Intermediate- Deformer 79.6$\pm$0.8 63.6±$\pm 0.3$ 23.0 35.7 68.6 76.4 4.39x 125x
Latency Sum-of-max 80.7±$\pm 0.2$ 58.8±$\pm 1.0$ 22.8 35.4 66.9 75.5 5.20x -
- w/ offline indexing 80.7±$\pm 0.2$ 58.8±$\pm 1.0$ 22.8 35.4 66.9 75.5 1.05x 125x
Low-Latency Bi-encoder 77.1 52.9 22.9 35.3 67.8 75.1 1x 1x
Poly-encoder 80.2±$\pm 0.1$ 57.6±$\pm 0.6$ 23.5 35.8 68.6 76.3 1.01x 1.0x
MixEncoder 75.4±$\pm 1.4$ 57.6±$\pm 0.3$ 20.7 32.5 68.2 75.8 1.12x 1.0x
CMC (Ours) 80.9±$\pm 0.1$ 59.2±$\pm 0.3$ 23.9 35.9 68.0 75.7 1.17x 1.0x
Table 3: Reranking Performance on four datasets with three downstream tasks: Entity Linking (Wikipedia-KB based datasets Hoffart et al. (2011); Guo and Barbosa (2018); Cucerzan (2007), ZeSHEL Logeswaran et al. (2019), Passage Ranking (MS MARCO Passage Ranking Bajaj et al. (2016), and Dialogue Ranking Gunasekara et al. (2019). The best result is denoted in bold and the second-best result is underlined. MRR stands for mean reciprocal rank and R@1 stands for recall at 1. In the entity linking datasets, the results are derived from averages across five random seeds. To show the computing resources required for the reranking process, we define reranking latency in terms of relative latency and additional memory usage compared to bi-encoders. indicates that the results are sourced from Wu et al. (2020); Yang et al. (2023); Yadav et al. (2022), respectively.
4.4 CMC as a Reranker
Baselines
In Table 3, baseline models are categorized into high-latency, intermediate-latency, and low-latency models. We adopt cross-encoders as our primary baseline for the high-latency approach. For the intermediate-latency models, we include Deformer and Sum-of-max, which utilize all vector embeddings to represent candidate information. For the low-latency models, we include the Bi-encoder, Poly-encoder, and Mixencoder, all of which require a single vector embedding for representation and have a serving time similar to that of the Bi-encoder. In this context, CMC is classified as a low-latency method because it requires a single embedding for the candidate and takes 1.17x serving time of the Bi-encoder.
Comparison with High-latency Models
Given the critical importance of computational resources and serving time in applications, CMC can be a highly practical alternative to cross-encoders. CMC shows 11.02x speedup and comparable prediction performance. CMC outperforms the cross-encoder in the Wikipedia entity linking (+0.7 accuracy points) and DSTC7 dataset (+3.3 MRR points) and presents a competitive performance in MS MARCO and ZeSHEL dataset, achieving the second- or third-best prediction performance. This comparison in performance underscores the significant role of the self-attention layer in CMC, suggesting that it effectively replaces the token-by-token interaction in cross-encoders with a balance between accuracy and computational efficiency.
Comparison with Intermediate-latency Models
When compared with intermediate-latency models such as Deformer and Sum-of-max, CMC demonstrates its capability not just in memory efficiency but also in maintaining competitive performance. CMC mostly surpasses these models in entity linking and passage ranking tasks. Also, CMC offers significant improvements in computational speed over Deformer (1.17x vs. 4.39x) and Sum-of-max without caching (1.17x vs. 5.20x).
For Sum-of-max with caching, it poses a critical issue in memory efficiency, requiring an index size 125 times larger than that of CMC. This gap becomes critical in terms of memory efficiency and scalability when managing extensive document collections. For instance, Sum-of-max and deformer require 3.2 TB to store just 5M entity descriptions from Wikipedia with 128 tokens, while CMC requires 23GB to store the same entities. As the knowledge base and document collections continuously change and grow, the operational feasibility of these systems quickly becomes impractical. Furthermore, the need to frequently update the embeddings to integrate the latest advances from pre-trained language models adds another layer of complexity and resource demand.
This analysis suggests that CMC’s single-vector approach is not only significantly faster under practical resource constraints but also demonstrates a comparable ability to represent candidate information with less information, occasionally surpassing more complex methods.
Comparison with Low-latency Models
In comparison with low-latency models, CMC matches or surpasses the performance of other baselines like Bi-encoder, Poly-encoder, and Mixencoder across different datasets. When compared with bi-encoders, the architectural distinctions contribute to CMC’s enhanced performance across every dataset. It shows the effectiveness of the CMC’s self-attention layer over the simple dot product. When evaluated against the Poly-encoder, CMC falls short on conversational datasets, as reported by Humeau et al. (2019). However, it outperforms Poly-encoder on other datasets not covered in the original baseline paper, which shows the robustness of CMC in handling more complex tasks than dialogue ranking. Similarly, CMC outperforms MixEncoder in entity linking and passage ranking.
4.5 Ablation Study
In Table 4, we explore the impact of different initialization checkpoints for sentence encoder on ZeSHEL reranking performance. We considered three sources for initialization: pre-trained BERT Devlin et al. (2018), ZeSHEL-fine-tuned bi-encoders Yadav et al. (2022), and MS-MARCO-fine-tuned-Cocondenser Guo and Barbosa (2018). For CMC, we find that initializing with a fine-tuned model significantly enhances performance. In contrast, Cocondenser leads to a performance decline in bi-encoders, highlighting CMC’s ability to facilitate domain adaptation of the sentence encoder. In poly-encoder and mixencoder, a similar trend to CMC is found, which can be attributed to their use of a single vector embedding from the candidate. On the other hand, Sum-of-max and Deformer demonstrate fairly consistent performance across various sentence encoders, likely because they use entire token embeddings, making them less sensitive to initialization choices.
5 Conclusion
In this paper, we present Comparing Multiple Candidates (CMC) which offers a novel approach to retrieve and rerank framework, addressing key issues in scalability and runtime efficiency. By utilizing language models for independent encoding and leveraging the self-attention layer, CMC achieves a balance of speed and effectiveness. Its ability to pre-compute candidate representations offline significantly reduces latency, making it a practical solution for enhancing end-to-end performance. Extensive experimentation validates CMC’s effectiveness, marking it as a promising advancement in the field of neural retrieval and reranking.
Bi-Encoder Initialization ZeSHEL
Methods
Vanilla
BERT
Fine-tuned
w/ ZeSHEL
Fine-tuned
w/ MS MARCO
Valid Test
Deformer $\checkmark$ 65.40 63.58
$\checkmark$ 64.42 62.43
$\checkmark$ 57.01 57.46
Sum-of-max $\checkmark$ 59.57 58.37
$\checkmark$ 58.77 57.65
$\checkmark$ 59.15 58.79
Bi-encoder $\checkmark$ 55.45 52.94
$\checkmark$ 49.32 44.01
Poly-encoder $\checkmark$ 53.37 52.49
$\checkmark$ 55.75 54.22
$\checkmark$ 57.41 58.22
MixEncoder $\checkmark$ 58.63 57.92
$\checkmark$ 58.52 57.68
$\checkmark$ 58.52 57.70
CMC(Ours) $\checkmark$ 56.15 55.34
$\checkmark$ 58.04 56.20
$\checkmark$ 60.05 59.23
Table 4: Comparison of unnormalized accuracy (U.Acc.) on ZeSHEL dataset over different initialization checkpoints (BERT Devlin et al. (2018), Bi-encoder Yadav et al. (2022) fine-tuned for ZeSHEL, and Cocondenser Guo et al. (2020)) fine-tuned for MS MARCO. We denote the best case for each method as bold.
Appendix A Limitations
In the future, we plan to test the CMC’s performance with over 1000 candidates using batch processing. This is because it has not yet been extensively researched whether CMC can effectively retrieve from a large collection, e.g. a collection comprising more than 1 million candidates. Furthermore, we plan to tackle the issue that arises from the concurrent operation of both a bi-encoder and CMC index, which currently requires double the index size. This is a consequence of running two separate encoder models in parallel. To address this, we will investigate various data compression techniques aimed at reducing the space footprint, thereby enhancing the practicality and efficiency of running both the Bi-encoder and CMC simultaneously.
Appendix B Potential Risks
This research examines methods to accelerate the two-stage retrieval and reranking process using efficient and effective CMC. While the proposed CMCmight exhibit specific biases and error patterns, we do not address these biases in this study. It remains uncertain how these biases might affect our predictions, an issue we plan to explore in future research.
Appendix C Detailed Information of Datasets
Wikipedia Entity Linking
For standard entity linking, we use AIDA-CoNLL dataset Hoffart et al. (2011) for in-domain evaluation, and WNED-CWEB Guo and Barbosa (2018) and MSNBC Cucerzan (2007) datasets for out-of-domain evaluation. These datasets share the same Wikipedia entity linking set. For comparison with the baseline results from Wu et al. (2020), we employ the 2019 English Wikipedia dump, containing 5.9M entities. We employed a bi-encoder as an initial retriever that yields an average unnormalized accuracy of 77.09 and a recall@10 of 89.21. Unnormalized accuracy is measured for each dataset and macro-averaged for test sets.
Regarding the license for each dataset, AIDA-CoNLL dataset is licensed under a Creative Commons Attribution-ShareAlike 3.0 Unported License. We are not able to find any license information about WNED-CWEB and MSNBC datasets.
Zero-shot Entity Linking (ZeSHEL)
ZeSHEL Logeswaran et al. (2019) contains mutually exclusive entity sets between training and test data. The dataset comprises context sentences (queries) each containing a mention linked to a corresponding gold entity description within Wikia knowledge base. The entity domain, also called “world”, varying from 10K to 100K entities, is unique to each domain, testing the model’s ability to generalize to new entities. We employed a bi-encoder from Yadav et al. (2022) whose recall@64 is 87.95. On top of these candidate sets, we report macro-averaged unnormalized accuracy, which is calculated for those mention sets that are successfully retrieved by the retriever and macro-averaged across a set of entity domains. For statistics of entity linking datasets, see Table 5. ZeSHEL is licensed under the Creative Commons Attribution-Share Alike License (CC-BY-SA).
The predominant approach for reranking in ZeSHEL dataset is based on top-64 candidate sets from official BM25 Logeswaran et al. (2019) or bi-encoder Wu et al. (2020); Yadav et al. (2022). On top of these candidate sets, we report macro-averaged normalized accuracy, which is calculated for those mention sets that are successfully retrieved by the retriever and macro-averaged across a set of entity domains.
Dataset # of Mentions # of Entities
AIDA Train 18848 5903530
Valid (A) 4791
Valid (B) 4485
MSNBC 656
WNED-WIKI 6821
ZeSHEL Train 49275 332632
Valid 10000 89549
Test 10000 70140
Table 5: Staistics of Entity Linking datasets.
MS MARCO
We use a popular passage ranking dataset MS MARCO which consists of 8.8 million web page passages. MS MARCO originates from Bing’s question-answering dataset with pairs of queries and passages, the latter marked as relevant if it includes the answer. Each query is associated with one or more relevant documents, but the dataset does not explicitly denote irrelevant ones, leading to the potential risk of false negatives. For evaluation, models are fine-tuned with approximately 500K training queries, and MRR@10, Recall@1 are used as a metric. To compare our model with other baselines, we employed Anserini’s BM25 Nogueira et al. (2019b). The dataset is licensed under Creative Commons Attribution 4.0 International.
DSTC 7 Challenge (Track 1)
For conversation ranking datasets, we involve The DSTC7 challenge (Track 1) Yoshino et al. (2019) . DSTC 7 involves dialogues taken from Ubuntu chat records, in which one participant seeks technical assistance for diverse Ubuntu-related issues. For these datasets, an official candidate set which includes gold is provided. For details for MS MARCO and DSTC 7 Challenge, see Table 6
Datasets Train Valid Test # of Candidates
per Query
MS MARCO 498970 6898 6837 1000
DSTC 7 100000 10000 5000 100
Table 6: Statistics of MS MARCO & Conversation Ranking Datasets.
Entity Linking Passage Ranking Dialogue Ranking
AIDA-train ZeSHEL MS MARCO DSTC7
Sentence Encoder Intialization BERT-large BERT-base BERT-base BERT-base
max. query length 32 128 32 512
max. document length 128 128 128 512
learning rate {1e-5,5e-6,2e-6} {1e-5,2e-5,5e-5} {1e-5,5e-6,2e-6} {1e-5,2e-5,5e-5}
batch size 4 4 8 8
hard negatives ratio 0.5 0.5 1 -
# of negatives 63 63 63 7
training epochs 4 5 3 10
Table 7: Hyperparameters for each dataset. We perform a grid search on learning rate and the best-performing learning rate is indicated as bold.
Appendix D Training Details
Negative Sampling
Most of previous studies that train reranker Wu et al. (2020); Xu et al. (2023) employ a fixed set of top-$k$ candidates from the retriever. In contrast, our approach adopts hard negative sampling, a technique derived from studies focused on training retrievers Zhang and Stratos (2021). Some negative candidates are sampled based on the retriever’s scoring for query-candidate pair $(q,c_{q,j})$:
$\forall j\in\{1,\ldots,K\}\setminus\{\text{gold index}\},$ (7)
$\quad\tilde{c}_{q,j}\sim\frac{\exp(s_{\text{retriever}}(q,\tilde{c}_{q,j}))}{\sum_{\begin{subarray}{c}k=1\\ k\neq\text{gold index}\end{subarray}}^{K}\exp(s_{\text{retriever}}(q,\tilde{c}_{q,k}))}$
To provide competitive and diverse negatives for the reranker, pof the negatives are fixed as the top-$k$ negatives, while the others are sampled following the score distribution.
As detailed in Table 7, we implement a hard negative mining strategy for training CMC and comparable baseline methods. Specifically, for the MS MARCO dataset, hard negatives are defined as the top 63 negatives derived from the CoCondenser model, as outlined in Gao and Callan (2022). In the case of entity linking datasets, we adhere to the approach established by Zhang and Stratos (2021), where hard negatives are selected from the top 1024 candidates generated by a bi-encoder. Meanwhile, for dialogue ranking datasets, we do not employ hard negative mining, owing to the absence of candidate pool within these datasets.
Sentence Encoder Initialization
The initial starting point for both the query and candidate encoders can significantly impact performance. The sentence encoders for late interaction models including CMC are initialized using either vanilla huggingface BERT Devlin et al. (2018) or other BERT-based, fine-tuned models. These models include those fine-tuned on the Wikipedia dataset (BLINK-biencoder; Wu et al. (2020)) or MS MARCO (Cocondenser; Gao and Callan (2022)). As cross-encoder is the only model without sentence encoder, we initialize cross-encoder only using vanilla BERT.
We initialize the sentence encoder for CMC and other baselines using (1) vanilla BERT and (2) the BLINK bi-encoder for Wikipedia entity linking datasets, and the MS-MARCO fine-tuned Cocondenser for other datasets. After conducting experiments with both starting points, we selected the best result among them. If more favorable results are found from prior works that conduct reranking over the same candidates, we sourced the numbers from these works.
Optimization
Our model employs multi-class cross-entropy as the loss function, regularized by Kullback-Leibler (KL) divergence between the reranker’s scores and the retriever’s scores. The loss function is formulated as follows:
$\begin{split}\mathcal{L}(q,\tilde{C}_{q})=-\lambda_{1}\sum_{i=1}^{K}y_{i}\log(p_{i})+\lambda_{2}\sum_{i=1}^{K}p_{i}\log\left(\frac{p_{i}}{r_{i}}\right)\\ \end{split}$ (8)
For the query $q$, $y_{i}$ represents the ground truth label for each candidate $\tilde{c}_{q,i}$, $p_{i}$ is the predicted probability for candidate $\tilde{c}_{q,i}$ derived from the score function $s_{\theta}$, $r_{i}$ is the probability of the same candidate from the retriever’s distribution, and $\lambda_{1}$ and $\lambda_{2}$ are coefficients forming a convex combination of the two losses.
Extra Skip Connection
[…]
Test Valid
Method R@1 R@4 R@8 R@16 R@32 R@64 R@1 R@64
Bi-encoder 52.94 64.51 71.94 81.52 84.98 87.95 55.45 92.04
Bi + CMC(64) 59.22 77.69 82.45 85.46 87.28 87.95 60.27 92.04
Bi + CMC(128) 59.13 77.65 82.72 85.84 88.29 89.83 60.24 93.22
Bi + CMC(256) 59.13 77.6 82.86 86.21 88.96 90.93 60.13 93.63
Bi + CMC(512) 59.08 77.58 82.91 86.32 89.33 91.51 60.1 93.89
Table 8: Retrieval performance by the number of candidates from the initial retriever. The numbers in parentheses (e.g., 128 for cmc(128)) indicate the number of candidates which CMC compares, initially retrieved by the bi-encoder. The best result is denoted in bold and the second-best result is underlined.
Appendix E Additional Results and Analysis
E.1 Reranking Latency of cross-encoders and CMC
Figure 3: The relationship between the number of candidates and the corresponding time measurements in milliseconds for two different models: Cross-encoder (CE) and Comparing Multiple Candidates (CMC).
In Figure 3, we present the plot of runtime against the number of candidates. For CMC, the model can handle up to 16,384 candidates per query, which is comparable to the speed of cross-encoders for running 64 candidates. Running more than 128 and 16,384 candidates cause memory error on GPU for cross-encoders and CMC, respectively.
E.2 Effect of Number of Candidates on Retrieval Performance
In Table 8, we present detailed results of retrieval performance on varying numbers of candidates from the initial bi-encoder. Recall@k increased with a higher number of candidates. It indicates that CMC enables the retrieval of gold instances that could not be retrieved by a bi-encoder, which prevents error propagation from the retriever. It is also noteworthy that CMC, which was trained using 64 candidates, demonstrates the capacity to effectively process and infer from a larger candidate pool (256 and 512) while giving an increase in recall@64 from 82.45 to 82.91.
E.3 Detailed Information of Entity Linking Performance
In Table 9, we present detailed results from Wikipedia entity linking datasets. Also, in table 10, we present detailed results for each world in ZeSHEL test set.
Method Valid (A) Test (B) MSNBC*
WNED-
CWEB*
Average
High- Cross-encoder 82.12 80.27 85.09 68.25 77.87
Latency Cross-encoder 87.15 83.96 86.69 69.11 80.22
Intermediate- Sum-of-max 90.84 85.30 86.07 70.65 80.67
Latency Deformer 90.64 84.57 82.92 66.97 78.16
Low- Bi-encoder 81.45 79.51 84.28 67.47 77.09
Latency Poly-encoder 90.64 84.79 86.30 69.39 80.16
MixEncoder 89.92 82.69 78.24 64.00 76.27
CMC 91.16 85.03 87.35 70.34 80.91
Table 9: Unnormalized accuracy on Wikipedia entity linking dataset (AIDA Hoffart et al. (2011), MSNBC Cucerzan (2007), and WNED-CWEB Guo and Barbosa (2018)). Average means macro-averaged accuracy for three test sets. The best result is denoted in bold and the second best result is denoted as underlined. * is out of domain dataset. is our implementation.
Valid Test (By Worlds)
Method
Forgotten
Realms
Lego Star Trek Yugioh Avg.
High- Cross-encoder 67.41 80.83 67.81 64.23 50.62 65.87
Latency
Cross-encoder
(w/ CMC)
70.22 81.00 67.89 64.42 50.86 66.04
Intermediate- Sum-of-max 59.15 73.45 58.83 57.63 45.29 58.80
Latency Deformer 56.95 73.08 56.98 56.24 43.55 57.46
Low- Bi-encoder 55.45 68.42 51.29 52.66 39.42 52.95
Latency Poly-encoder 57.19 71.95 58.11 56.19 43.60 57.46
MixEncoder 58.64 73.17 56.29 56.99 43.01 57.36
CMC(Ours) 60.05 73.92 58.96 58.08 45.69 59.16
Table 10: Detailed Reranking Performance on Zero-shot Entity Linking (ZeSHEL) valid and test set Logeswaran et al. (2019). Macro-averaged unnormalized accuracy is measured for candidates from Bi-encoder Yadav et al. (2022).The best result is denoted in bold.
E.4 Ranking Performance on ZeSHEL BM25 candidate sets
In many previous works Wu et al. (2020); Xu et al. (2023), the performance of models over BM25 candidates Logeswaran et al. (2019) has been reported. In Table 11, we present the performance of CMC to illustrate its positioning within this research landscape.
Methods
Forgotten
Realms
Lego Star Trek Yugioh Macro Acc. Micro Acc.
Cross-encoder Wu et al. (2020) 87.20 75.26 79.61 69.56 77.90 77.07
ReS Xu et al. (2023) 88.10 78.44 81.69 75.84 81.02 80.40
ExtEnD De Cao et al. (2020) 79.62 65.20 73.21 60.01 69.51 68.57
GENRE De Cao et al. (2020) 55.20 42.71 55.76 34.68 47.09 47.06
Poly-encoder 78.90 64.47 71.05 56.25 67.67 66.81
Sum-of-max 83.20 68.17 73.14 64.00 72.12 71.15
Comparing Multiple Candidates (Ours) 83.20 70.63 75.75 64.83 73.35 72.41
Table 11: Test Normalized accuracy of CMC model over retrieved candidates from BM25. is reported from Xu et al. (2023). is our implementation.
E.5 Ablation Study on Training Strategies
In Table 12, we evaluated the impact of different training strategies on the CMC’s reranking performance on the ZeSHEL test set. The removal of extra skip connections results in only a slight decrease ranging from 0.03 to 0.39 points in normalized accuracy. Also, to examine the effects of a bi-encoder retriever, we remove regularization from the loss. It leads to a performance drop but still shows higher performance than ColBERT, the most powerful baseline in the low latency method. Lastly, we tried to find the influence of negative sampling by using fixed negatives instead of mixed negatives. The result shows a marginal decline in the test set, which might be due to the limited impact of random negatives in training CMC.
w/ bi-encoder retriever w/ BM25 retriever
Methods Valid Test Test
CMC 65.29 66.83 73.1
w/o extra skip connection 64.78 66.44 73.07
w/o regularization 64.45 66.31 72.94
fixed negative 65.32 66.46 72.97
Table 12: Normalized Accuracy on ZeSHEL test set for various training strategies
E.6 Reranking Performance of Cross-encoders for Various Number of Candidates
In Table 13, we evaluated the impact of the different number of candidates on the cross-encoder’s reranking performance on the ZeSHEL test set with a candidate set from the bi-encoder retriever. Even with a larger number of candidates, the unnormalized accuracy of the cross-encoder does not increase. Although the number of candidates from the bi-encoder increases from 64 to 512, recall@1 decreases by 0.01 points.
# of candidates
Recall@1
(Unnormalized Accuracy)
16 65.02
64 65.87
512 65.85
Table 13: Normalized Accuracy on ZeSHEL test set for various training strategies